# Tensor Product Attention Is All You Need

### v6

@@ Abstract
Scaling language models to handle longer input sequences typically necessitates large key-value (KV) caches, resulting in substantial memory overhead during inference. In this paper, we propose Tensor Product Attention (TPA), a novel attention mechanism that uses tensor decompositions to represent queries, keys, and values compactly, substantially shrinking the KV cache size at inference time. By factorizing these representations into contextual low-rank components and seamlessly integrating with Rotary Position Embedding (RoPE), TPA achieves improved model quality alongside memory efficiency. Based on TPA, we introduce the Tensor ProducT ATTenTion Transformer (T6), a new model architecture for sequence modeling. Through extensive empirical evaluation on language modeling tasks, we demonstrate that T6 surpasses or matches the performance of standard Transformer baselines, including Multi-Head Attention (MHA), Multi-Query Attention (MQA), Grouped-Query Attention (GQA), and Multi-Head Latent Attention (MLA) across various metrics, including perplexity and a range of established evaluation benchmarks. Notably, TPA’s memory efficiency and computational efficiency at the decoding stage enable processing longer sequences under fixed resource constraints, addressing a critical scalability challenge in modern language models. Project Page: https://github.com/tensorgi/TPA.
@@ 1 Introduction
Large language models (LLMs) have revolutionized natural language processing, demonstrating exceptional performance across tasks (brown2020language; chowdhery2022palm; touvron2023llama; bubeck2023sparks). As these models evolve, their ability to process longer contexts becomes increasingly important for sophisticated applications such as document analysis, complex reasoning, and code completion. However, managing longer sequences during inference poses significant computational and memory challenges, particularly due to the storage of key-value (KV) caches (zhang2023h2o; liu2024kivi). Because memory consumption grows linearly with sequence length, the maximum context window is limited by practical hardware constraints.
@@
A variety of solutions have been explored to address this memory bottleneck. Some approaches compress or selectively prune cached states through sparse attention patterns (child2019generating) or token eviction strategies (zhang2023h2o; xiao2023efficient; ribar2023sparq), though such methods risk discarding tokens that may later prove important. Other work proposes off-chip storage of key-value states (he2024fastdecode), at the expense of increased I/O latency. Attention variants like Multi-Query Attention (MQA) (shazeer2019fast) and Grouped-Query Attention (GQA) (ainslie2023gqa) reduce per-token cache requirements by sharing keys and values across heads, but often compromise flexibility or require significant architectural modifications. Meanwhile, low-rank weight factorization methods such as LoRA (hu2021lora) effectively reduce fine-tuning memory, yet do not address the KV cache overhead that dominates inference at runtime. The recently introduced Multi-Head Latent Attention (MLA) in Deepseek-V2 (liu2024deepseek) caches compressed key-value representations but encounters difficulties with efficient Rotary Position Embedding (RoPE) (su2024roformer) integration, necessitating additional position-encoded parameters per head.
To overcome the limitations of existing approaches, we introduce Tensor Product Attention (TPA), illustrated in Figure 1. TPA is a novel attention mechanism that employs tensor factorizations for queries (Q), keys (K), and values (V). By dynamically factorizing activations rather than static weights (as in LoRA), TPA constructs low-rank, contextual representations. This approach substantially reduces KV cache memory usage while offering improved representational capacity. In practice, TPA can decrease memory overhead by an order of magnitude compared to standard Multi-Head Attention (MHA), alongside achieving lower pretraining validation loss (perplexity) and better downstream performance. A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (su2024roformer) and any possible position encodings, enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (touvron2023llama), Qwen (bai2023qwen), and Gemma (team2024gemma).
@@
    We propose Tensor Product Attention (TPA), a mechanism that factorizes $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ activations using contextual tensor decompositions. This achieves a substantial reduction in inference-time KV cache size relative to standard attention mechanisms (vaswani2017attention), MHA, MQA, GQA, and MLA, while also improving performance. In addition, we analyze existing attention mechanisms and reveal that MHA, MQA, and GQA can be expressed as non-contextual variants of TPA.
@@
    We demonstrate that TPA integrates seamlessly with RoPE (su2024roformer) and any possible position encodings as well as output gate and KV shifting, facilitating its easy adoption in popular foundation model architectures like LLaMA, Gemma, and Qwen.
@@ 2 Background
In this section, we briefly review Scaled Dot-Product Attention, Multi-Head Attention (vaswani2017attention), and introduce key notations. Other attention mechanisms like Multi-Query Attention (MQA) (shazeer2019fast), Grouped Query Attention (GQA) (ainslie2023gqa), Multi-head Latent Attention (MLA) (liu2024deepseek; liu2024deepseekv3), and Rotary Position Embedding (RoPE) (su2024roformer) are further discussed in the Appendix F.
Notations. We use bold uppercase letters (e.g., $\mathbf{X}$, $\mathbf{Q}$) for matrices, bold lowercase (e.g., $\mathbf{a}$, $\mathbf{b}$) for vectors, and italic uppercase (e.g., $\bm{W}^{Q}_{i}$) for learnable parameter matrices. We denote by $[n]$ the set $\{1,\ldots,n\}$ for some positive integer $n$. We use $\top$ to denote the transpose of a vector or a matrix. Let $d_{\text{model}}$ be the embedding dimension, $h$ the number of attention heads, $d_{h}$ the dimension per head, $\mathbf{x}_{t}\in\mathbb{R}^{d}$ the input for the $t$-th token at a given attention layer, $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$ denotes the input embeddings for $T$ tokens, and $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ denote the queries, keys, and values of $h$ heads for $T$ tokens. With a little abuse of notation, $\mathbf{Q}_{i}$, $\mathbf{K}_{i}$, $\mathbf{V}_{i}\in\mathbb{R}^{T\times d_{h}}$ denote the $i$-th head of queries, keys, and values, and $\mathbf{Q}_{t}$, $\mathbf{K}_{t}$, $\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ denote the heads of the query, key, and value for $t$-th token. Throughout the paper, $\bm{W}^{Q},\bm{W}^{K},\bm{W}^{V}$ denote projection matrices for queries, keys, and values, respectively. In multi-head attention, each head is associated with its own set of $\bm{W}^{Q}_{i},\bm{W}^{K}_{i},\bm{W}^{V}_{i}$, and each has dimension $\bm{W}^{Q}_{i},\bm{W}^{K}_{i},\bm{W}^{V}_{i}\in\mathbb{R}^{\,d_{\text{model}}\times d_{k}}$, where $d_{k}$ is typically set to $d_{h}$, the dimension of each head. [^5] Often, $h\times d_{h}=d_{\text{model}}$, so each head has query/key/value dimension $d_{h}$. Similarly, we have an output projection matrix $\bm{W}^{O}\in\mathbb{R}^{(h\cdot d_{h})\times d_{\text{model}}}$. For methods like MQA and GQA, some of these projection matrices are shared or partially shared across heads, but their shapes remain consistent.
@@ 2.1 Scaled Dot-Product Attention
Scaled dot-product attention (vaswani2017attention) determines how to focus on different parts of an input sequence by comparing queries ($\mathbf{Q}$) and keys ($\mathbf{K}$). It produces a weighted combination of the values ($\mathbf{V}$). Formally, the attention output is:
$\operatorname{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{Softmax}\Bigl(\tfrac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\Bigr)\,\mathbf{V},$
where each of $\mathbf{Q},\mathbf{K},\mathbf{V}$ is an $(n\times d_{k})$ matrix for $n$ tokens and key dimension $d_{k}$.
@@ 2.2 Multi-Head Attention (MHA)
Multi-Head Attention (MHA) (vaswani2017attention) extends scaled dot-product attention by dividing the model’s internal representation into several heads. Each head learns different projections for queries, keys, and values, allowing the model to attend to different types of information from different representational subspaces. For each token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$, MHA computes each head $i$ as follows:
@@ 3.1 Tensor Factorization of Queries, Keys, and Values
Let $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$ for $t=1,\dots,T$ be the hidden-state vector corresponding to the $t$-th token in a sequence of length $T$. A typical multi-head attention block has $h$ heads, each of dimension $d_{h}$, satisfying $d_{\text{model}}=h\times d_{h}$. Standard attention projects the entire sequence into three tensors, $\mathbf{Q},\;\mathbf{K},\;\mathbf{V}\;\in\mathbb{R}^{T\times h\times d_{h}}$, where $\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ denote the slices for the $t$-th token.
Contextual Factorization. Instead of forming each head’s query, key, or value via a single linear map, TPA factorizes each $\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}$ into a sum of (contextual) tensor products whose ranks are $R_{q}$, $R_{k}$, and $R_{v}$, respectively and may differ. Specifically, for each token $t$, with a small abuse of notation, we define:
@@
where $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h}$$\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}}$. Hence, for queries, each tensor product $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\colon\mathbb{R}^{h}\times\mathbb{R}^{d_{h}}\to\mathbb{R}^{h\times d_{h}}$ (an outer product) contributes to the query slice $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$. Analogous definitions apply to the key slice $\mathbf{K}_{t}$ and value slice $\mathbf{V}_{t}$.
@@
Parameter Initialization. We use Xavier initialization (glorot2010understanding) for the factor weight matrices; details are in the Appendix G.
@@ 3.2 RoPE Compatibility and Acceleration
$\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{t})\;\longleftarrow\;\operatorname{RoPE}_{t}\bigl(\mathbf{B}_{K}(\mathbf{x}_{t})\bigr),$ (3.5)
@@
$\widetilde{\mathbf{K}}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{(r)}(\mathbf{x}_{t})\otimes\operatorname{RoPE}_{t}\bigl(\mathbf{b}^{K}_{(r)}(\mathbf{x}_{t})\bigr)=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{t})^{\top}\operatorname{RoPE}_{t}\bigl(\mathbf{B}_{K}(\mathbf{x}_{t})\bigr).$
Here, $\operatorname{RoPE}_{t}$ is applied to each row of $\mathbf{B}_{K}(\mathbf{x}_{t})$ (i.e., to each $\mathbf{b}^{K}_{(r)}(\mathbf{x}_{t})$ vector). Thus, each $\mathbf{K}_{t}$ is effectively rotated before caching. This removes the need for explicit rotation at decoding time, accelerating autoregressive inference. Depending on hardware and performance requirements, different RoPE integration strategies can be adopted for training and inference.
@@
$\operatorname{RoPE}({\mathbf{Q}}_{t})=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\,\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}),$ (3.6)
where $\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t})=\operatorname{RoPE}_{t}\bigl(\mathbf{B}_{Q}(\mathbf{x}_{t})\bigr)$ (RoPE applied row-wise to $\mathbf{B}_{Q}(\mathbf{x}_{t})$). Furthermore, let $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ be factorized by TPA. Let $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}(\mathbf{Q}_{t})$ and $\widetilde{\mathbf{K}}_{s}=\operatorname{RoPE}_{s}(\mathbf{K}_{s})$ be their RoPE-transformed versions. The relative positional encoding property of RoPE is preserved:
where $\operatorname{RoPE}_{t-s}$ denotes applying RoPE with relative position tst-s. Focusing on individual heads $i$, if $\mathbf{q}_{t,i}$ and $\mathbf{k}_{s,i}$ are the $i$-th head vectors from $\mathbf{Q}_{t}$ and $\mathbf{K}_{s}$ respectively (as column vectors of dimension $d_{h}$), and $\widetilde{\mathbf{q}}_{t,i}=\operatorname{RoPE}(\mathbf{q}_{t,i},t)$ and $\widetilde{\mathbf{k}}_{s,i}=\operatorname{RoPE}(\mathbf{k}_{s,i},s)$, then the dot product for attention scores satisfies:
$\widetilde{\mathbf{q}}_{t,i}^{\top}\,\widetilde{\mathbf{k}}_{s,i}=\mathbf{q}_{t,i}^{\top}\operatorname{RoPE}(\cdot,t-s)\mathbf{k}_{s,i}.$

### v7

@@ Abstract
Scaling language models to handle longer input sequences typically necessitates large key-value (KV) caches, resulting in substantial memory overhead during inference. In this paper, we propose Tensor Product Attention (TPA), a novel attention mechanism that uses tensor decompositions to represent queries, keys, and values compactly, substantially shrinking the KV cache size at inference time. By factorizing these representations into contextual low-rank components and seamlessly integrating with RoPE and any possible position encoding mechanisms, TPA achieves improved model quality alongside memory efficiency. Based on TPA, we introduce the Tensor ProducT ATTenTion Transformer (T6), a new model architecture for sequence modeling. Through extensive empirical evaluation on language modeling tasks, we demonstrate that T6 surpasses or matches the performance of standard Transformer baselines, including Multi-Head Attention (MHA), Multi-Query Attention (MQA), Grouped-Query Attention (GQA), and Multi-Head Latent Attention (MLA) across various metrics, including perplexity and a range of established evaluation benchmarks. Notably, TPA’s memory efficiency and computational efficiency at the decoding stage enable processing longer sequences under fixed resource constraints, addressing a critical scalability challenge in modern language models. Project Page: https://github.com/tensorgi/TPA.
@@ 1 Introduction
Large language models (LLMs) have revolutionized natural language processing, demonstrating exceptional performance across tasks (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023; Bubeck et al., 2023). As these models evolve, their ability to process longer contexts becomes increasingly important for sophisticated applications such as document analysis, complex reasoning, and code completion. However, managing longer sequences during inference poses significant computational and memory challenges, particularly due to the storage of key-value (KV) caches (Zhang et al., 2023c; Liu et al., 2024c). Because memory consumption grows linearly with sequence length, the maximum context window is limited by practical hardware constraints.
@@
A variety of solutions have been explored to address this memory bottleneck. Some approaches compress or selectively prune cached states through sparse attention patterns (Child et al., 2019) or token eviction strategies (Zhang et al., 2023c; Xiao et al., 2024; Ribar et al., 2024), though such methods risk discarding tokens that may later prove important. Other work proposes off-chip storage of key-value states (He and Zhai, 2024), at the expense of increased I/O latency. Attention variants like Multi-Query Attention (MQA) (Shazeer, 2019) and Grouped-Query Attention (GQA) (Ainslie et al., 2023) reduce per-token cache requirements by sharing keys and values across heads, but often compromise flexibility or require significant architectural modifications. Meanwhile, low-rank weight factorization methods such as LoRA (Hu et al., 2022) effectively reduce fine-tuning memory, yet do not address the KV cache overhead that dominates inference at runtime. The recently introduced Multi-Head Latent Attention (MLA) in Deepseek-V2 (Liu et al., 2024a) caches compressed key-value representations but encounters difficulties with efficient Rotary Position Embedding (RoPE) (Su et al., 2024) integration, necessitating additional position-encoded parameters per head.
To overcome the limitations of existing approaches, we introduce Tensor Product Attention (TPA), illustrated in Figure 1. TPA is a novel attention mechanism that employs tensor factorizations for queries (Q), keys (K), and values (V). By dynamically factorizing activations rather than static weights (as in LoRA), TPA constructs low-rank, contextual representations. This approach substantially reduces KV cache memory usage while offering improved representational capacity. In practice, TPA can decrease memory overhead by an order of magnitude compared to standard Multi-Head Attention (MHA), alongside achieving lower pretraining validation loss (perplexity) and better downstream performance. A key advantage of TPA is its native compatibility with rotary positional embeddings (RoPE) (Su et al., 2024) and any possible position encodings, enabling a straightforward drop-in replacement for multi-head attention (MHA) layers in modern LLM architectures such as LLaMA (Touvron et al., 2023), Qwen (Bai et al., 2023), and Gemma (Team et al., 2024).
@@
    We propose Tensor Product Attention (TPA), a mechanism that factorizes $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ activations using contextual tensor decompositions. This achieves a substantial reduction in inference-time KV cache size relative to standard attention mechanisms (Vaswani et al., 2017), MHA, MQA, GQA, and MLA, while also improving performance. In addition, we analyze existing attention mechanisms and reveal that MHA, MQA, and GQA can be expressed as non-contextual variants of TPA.
@@
    We demonstrate that TPA integrates seamlessly with RoPE (Su et al., 2024) and any possible position encodings as well as output gate and KV shifting, facilitating its easy adoption in popular foundation model architectures like LLaMA, Gemma, and Qwen.
@@ 2 Background
In this section, we briefly review Scaled Dot-Product Attention, Multi-Head Attention (Vaswani et al., 2017), and introduce key notations. Other attention mechanisms like Multi-Query Attention (MQA) (Shazeer, 2019), Grouped Query Attention (GQA) (Ainslie et al., 2023), Multi-head Latent Attention (MLA) (Liu et al., 2024a, b), and Rotary Position Embedding (RoPE) (Su et al., 2024) are further discussed in the Appendix F.
Notations. We use bold uppercase letters (e.g., $\mathbf{X}$, $\mathbf{Q}$) for matrices, bold lowercase (e.g., $\mathbf{a}$, $\mathbf{b}$) for vectors, and italic uppercase (e.g., $\bm{W}^{Q}_{i}$) for learnable parameter matrices. We denote by $[n]$ the set $\{1,\ldots,n\}$ for some positive integer $n$. We use $\top$ to denote the transpose of a vector or a matrix. Let $d_{\text{model}}$ be the embedding dimension, $h$ the number of attention heads, $d_{h}$ the dimension per head, $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$ the input for the $t$-th token at a given attention layer, $\mathbf{X}\in\mathbb{R}^{T\times d_{\text{model}}}$ denotes the input embeddings for $T$ tokens, and $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}\in\mathbb{R}^{T\times h\times d_{h}}$ denote the queries, keys, and values of $h$ heads for $T$ tokens. With a little abuse of notation, $\mathbf{Q}_{i}$, $\mathbf{K}_{i}$, $\mathbf{V}_{i}\in\mathbb{R}^{T\times d_{h}}$ denote the $i$-th head of queries, keys, and values, and $\mathbf{Q}_{t}$, $\mathbf{K}_{t}$, $\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ denote the heads of the query, key, and value for $t$-th token. Throughout the paper, $\bm{W}^{Q},\bm{W}^{K},\bm{W}^{V}$ denote projection matrices for queries, keys, and values, respectively. In multi-head attention, each head is associated with its own set of $\bm{W}^{Q}_{i},\bm{W}^{K}_{i},\bm{W}^{V}_{i}$, and each has dimension $\bm{W}^{Q}_{i},\bm{W}^{K}_{i},\bm{W}^{V}_{i}\in\mathbb{R}^{\,d_{\text{model}}\times d_{h}}$. [^5] Often, $h\times d_{h}=d_{\text{model}}$, so each head has query/key/value dimension $d_{h}$. Similarly, we have an output projection matrix $\bm{W}^{O}\in\mathbb{R}^{(h\cdot d_{h})\times d_{\text{model}}}$.
@@ 2.1 Scaled Dot-Product Attention
Scaled dot-product attention (Vaswani et al., 2017) determines how to focus on different parts of an input sequence by comparing queries ($\mathbf{Q}$) and keys ($\mathbf{K}$). It produces a weighted combination of the values ($\mathbf{V}$). Formally, the attention output is:
$\operatorname{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\operatorname{Softmax}\Bigl(\tfrac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{h}}}\Bigr)\,\mathbf{V},$
where $\mathbf{Q}\in\mathbb{R}^{n\times d_{h}}$, $\mathbf{K}\in\mathbb{R}^{n\times d_{h}}$, and $\mathbf{V}\in\mathbb{R}^{n\times d_{v}}$ for $n$ tokens. The softmax is applied row-wise over the $n$ keys for each query.
@@ 2.2 Multi-Head Attention (MHA)
Multi-Head Attention (MHA) (Vaswani et al., 2017) extends scaled dot-product attention by dividing the model’s internal representation into several heads. Each head learns different projections for queries, keys, and values, allowing the model to attend to different types of information from different representational subspaces. For each token embedding $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{model}}}$, MHA computes each head $i$ as follows:
@@ 3.1 Tensor Factorization of Queries, Keys, and Values
Let $d_{\text{attn}}:=h\,d_{h}$ denote the total attention projection dimension. Typically one sets $d_{\text{attn}}=d_{\text{model}}$, but this is not required: when $d_{\text{attn}}\neq d_{\text{model}}$, the projection matrices $\bm{W}^{Q},\bm{W}^{K},\bm{W}^{V}$ map from $\mathbb{R}^{d_{\text{model}}}$ into $\mathbb{R}^{d_{\text{attn}}}$ and $\bm{W}^{O}$ maps $\mathbb{R}^{d_{\text{attn}}}$ back to $\mathbb{R}^{d_{\text{model}}}$. Standard attention projects the entire sequence into three tensors, $\mathbf{Q},\;\mathbf{K},\;\mathbf{V}\;\in\mathbb{R}^{T\times h\times d_{h}}$, where $\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}\in\mathbb{R}^{h\times d_{h}}$ denote the slices for the $t$-th token.
Contextual Factorization. Instead of forming each head’s query, key, or value via a single linear map, TPA factorizes each $\mathbf{Q}_{t},\mathbf{K}_{t},\mathbf{V}_{t}$ into a sum of (contextual) tensor products whose ranks are $R_{Q}$, $R_{K}$, and $R_{V}$, respectively, and may differ. Specifically, for each token $t$, with a small abuse of notation, we define:
@@
where $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{a}^{K}_{r}(\mathbf{x}_{t}),\mathbf{a}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h}$$\mathbf{b}^{Q}_{r}(\mathbf{x}_{t}),\mathbf{b}^{K}_{r}(\mathbf{x}_{t}),\mathbf{b}^{V}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{h}}$. Hence, for queries, each tensor product $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\colon\mathbb{R}^{h}\times\mathbb{R}^{d_{h}}\to\mathbb{R}^{h\times d_{h}}$ contributes to the query slice $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$. Analogous definitions apply to the key slice $\mathbf{K}_{t}$ and value slice $\mathbf{V}_{t}$.
@@
Parameter Initialization. We use Xavier initialization (Glorot and Bengio, 2010) for the factor weight matrices; details are in the Appendix G.
@@ 3.2 RoPE Compatibility and Acceleration
$\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{t})\;:=\;\operatorname{RoPE}_{t}\bigl(\mathbf{B}_{K}(\mathbf{x}_{t})\bigr)=\mathbf{B}_{K}(\mathbf{x}_{t})\mathbf{T}_{t},$ (3.5)
@@
$\widetilde{\mathbf{K}}_{t}=\frac{1}{R_{K}}\sum_{r=1}^{R_{K}}\mathbf{a}^{K}_{r}(\mathbf{x}_{t})\otimes\operatorname{RoPE}_{t}\bigl(\mathbf{b}^{K}_{r}(\mathbf{x}_{t})\bigr)=\frac{1}{R_{K}}\mathbf{A}_{K}(\mathbf{x}_{t})^{\top}\widetilde{\mathbf{B}}_{K}(\mathbf{x}_{t}).$
Here, $\operatorname{RoPE}_{t}$ is applied to each row of $\mathbf{B}_{K}(\mathbf{x}_{t})$ (i.e., to each $\mathbf{b}^{K}_{r}(\mathbf{x}_{t})$ vector). Thus, each cached key factor corresponds to a RoPE-rotated key slice. This removes the need to rotate cached keys at decoding time; the current-step query (which is not cached) can still be rotated on the fly at negligible cost. Depending on hardware and performance requirements, different RoPE integration strategies can be adopted for training and inference.
@@
$\operatorname{RoPE}_{t}(\mathbf{Q}_{t})=\mathbf{Q}_{t}\mathbf{T}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\,\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}),$ (3.6)
where $\widetilde{\mathbf{B}}_{Q}(\mathbf{x}_{t}):=\mathbf{B}_{Q}(\mathbf{x}_{t})\mathbf{T}_{t}=\operatorname{RoPE}_{t}\bigl(\mathbf{B}_{Q}(\mathbf{x}_{t})\bigr)$ (RoPE applied row-wise to $\mathbf{B}_{Q}(\mathbf{x}_{t})$). Furthermore, let $\widetilde{\mathbf{Q}}_{t}=\operatorname{RoPE}_{t}(\mathbf{Q}_{t})=\mathbf{Q}_{t}\mathbf{T}_{t}$ and $\widetilde{\mathbf{K}}_{s}=\operatorname{RoPE}_{s}(\mathbf{K}_{s})=\mathbf{K}_{s}\mathbf{T}_{s}$ be the RoPE-transformed query/key slices. Then RoPE’s standard relative-position identity is preserved:
$\widetilde{\mathbf{Q}}_{t}\,\widetilde{\mathbf{K}}_{s}^{\top}\;=\;\mathbf{Q}_{t}\mathbf{T}_{t-s}\mathbf{K}_{s}^{\top},\qquad\text{equivalently}\qquad\operatorname{RoPE}_{t-s}(\mathbf{Q}_{t})\,\mathbf{K}_{s}^{\top}\;=\;\widetilde{\mathbf{Q}}_{t}\,\widetilde{\mathbf{K}}_{s}^{\top},$
where $\mathbf{T}_{t-s}:=\mathbf{T}_{t}\mathbf{T}_{s}^{\top}$. In particular, for any head $i$ (the $i$-th row), if $\mathbf{q}_{t,i},\mathbf{k}_{s,i}\in\mathbb{R}^{1\times d_{h}}$ and $\widetilde{\mathbf{q}}_{t,i}=\mathbf{q}_{t,i}\mathbf{T}_{t}$, $\widetilde{\mathbf{k}}_{s,i}=\mathbf{k}_{s,i}\mathbf{T}_{s}$, then $\widetilde{\mathbf{q}}_{t,i}\,\widetilde{\mathbf{k}}_{s,i}^{\top}\;=\;\mathbf{q}_{t,i}\mathbf{T}_{t-s}\mathbf{k}_{s,i}^{\top}$.